Two fair dice are tossed simultaneously and the bigger of the two numbers obtained is recorded. If A represent the event 'bigger number is 4' and B represents the event 'bigger number is odd', find:

a.) P(A)

b.) P(A U B)


1
Expert's answer
2021-05-17T18:50:19-0400

"A\to" Bigger number is 4


(i) Sample sapce for A is {"(1,4),(2,4),(3,4),(4,1),(4,2),(4,3)" }


So, "P(A)=\dfrac{6}{36}=0.166"


(ii) "B\to" Bigger number is odd


So, "P(B)=\dfrac{10}{36}=0.27"


"P(A\cap B)=0"


Hence, "P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=0.166+0.27-0\\=0.436"